Alexander Perez and Alexandra Williams are senior vice presidents of Columbia Group Health Cooperative (CGHC) and co-directors of the organization's pension fund management division. The unions that represent the CGHC staff have requested an investment seminar so that they better understand the decisions being made on behalf of their members. Alexander and Alexandra, who will make the actual presentation, have asked you to help them. To illustrate the common stock valuation process, Alexander and Alexandra have asked you to analyze the Temp Force Company, an employment agency that supplies word processor operators and computer programmers to businesses with temporarily heavy workloads. You are to answer the following questions.                                                                                               

a. Assume that Temp Force has a beta coefficient of 1.4, that the risk-free rate (the yield on T-bonds) is 6 percent, and that the market risk premium is 5 percent. What is the required rate of return on the firm’s stock?         

b. Assume that Temp Force is a constant growth company whose last dividend (D0, which was paid yesterday) was $1.80, and whose dividend is expected to grow indefinitely at a 9 percent rate. What is the firm’s expected dividend stream over the next three years?               

c. What is the firm’s expected current stock price?                   

d. What is the stock's expected value one year from now?

e. What are the expected dividend yield, capital gains yield, and total return during the first year?

f. What would the expected current stock price be if its dividends were expected to have zero growth?
Answer #1

a. Given beta coefficient Beta = 1.4

Given risk-free rate Rf = 6%

Given market risk premium Rm - Rf= 5%

Hence Required rate of return = Rf + Beta*( Rm - Rf)

= 6 + 1.4(5) = 12%

b. Given D0 = $1.80

Given Growth rate = 9%

Year 1 Year 2 Year 3
Dividend 1.962 2.13 2.33

c. From the Dividend discount model -

(Price of stock )0 = D1/(r-g)

= 1.962/(.12-.09) = $65.4

d. Stocks value 1 year from now

P1 = D2/ (r-g) = 2.13/(.12-.09) = $71

e.If dividends were not expected to grow then -

D0 = D1 /(r-g) = 1.8(1)/(.12-.09) = $60

i.e. as if the stock had a 0 growth rate